# On $3$-manifolds admitting co-orientable taut foliations, but none with vanishing Euler class

### v2

@@ 3. The Euler class of the normal bundle of a Seifert fibration
Let $M$ be a closed, oriented Seifert fibred manifold whose base orbifold has an orientable underlying space. In this section, we compute the Euler class of the normal bundle of the Seifert fibration on $M$ and determine conditions for it to vanish (Theorem 1.11 and Corollary 1.12).
@@ 3.1. Seifert notation
    This description of the Seifert fibred manifold is not unique. More precisely, we have that $M\Bigl(g;\,\tfrac{a_{0}}{b_{0}},\ldots,\tfrac{a_{n}}{b_{n}}\Bigr)$ is homeomorphic to $M\Bigl(g;\,\tfrac{a^{\prime}_{0}}{b_{0}},\ldots,\tfrac{a^{\prime}_{n}}{b_{n}}\Bigr)$ by an orientation-preserving homeomorphism preserving fibres if and only if up to reordering,
@@
It follows from [13, Theorem 4] that the image of 1-1 under the transgression homomorphism $\delta$ is the Euler class $bu$ of the circle bundle. Hence $\mbox{kernel}(f^{*})$ is generated by $bu$. Then by (3.1.2), $e(\nu_{M})=0$ if and only $e(\tau_{B})=\pm\chi(B)u$ is a multiple of $bu$, which is condition (3.1.1).
@@
    From $\sigma$ one can define an obstruction $2$-cochain, denoted by $c_{\sigma}$, that represents the relative Euler class of the normal bundle $\nu_{M}$ in $H^{2}(M,\sqcup T_{i})$. By construction, the image of $c_{\sigma}]$ under $H^{2}(M,\sqcup T_{i})\to H^{2}(M)$ is the Euler class $e(\nu_{M})$. (See [14, Section 2] for instance.)
@@
4. Discrete representations of Fuchsian groups and the normal bundle of a Seifert fibration
@@ 4.2. The normal bundle of the Seifert fibring of $M$ and discrete representations
Now proceed as in the proof of [3, Lemma A.3], starting from its third paragraph, to show that the $\mathbb{H}^{2}$-bundle $M\times_{\psi}\mathbb{H}^{2}\to M$ is topologically equivalent to the normal bundle $\nu_{M}$ and conclude, as in [3, Proposition A4], that $e(\nu_{M})=e(\varphi)=e(\bar{\varphi}\circ\alpha)$. Thus
@@
It is well-known that $e(\bar{\varphi})$, respectively $e(\varphi)$, is the obstruction to lifting the representation $\bar{\varphi}:\Gamma\to PSL_{2}(\mathbb{R})$, respectively $\varphi:\pi_{1}(M)\to PSL_{2}(\mathbb{R})$, to $\widetilde{PSL}_{2}(\mathbb{R})$ [19, Lemma 2].
@@
(4.2.4) $\theta(e(E_{M}))=-l(b+\sum_{i=1}^{n}\frac{a_{i}}{b_{i}})=-le(M),$
@@
(4.2.5) $\theta(e(\bar{\varphi}))=l(n-\chi(|\mathcal{B}|)-\sum_{i}\frac{1}{b_{i}})=-l\chi(\mathcal{B})$
@@
We saw in (4.2.1) that that $e(\nu_{M})=\alpha^{*}(e(\bar{\varphi}))=e(\bar{\varphi}\circ\alpha)=e(\varphi)$, where $\varphi=\bar{\varphi}\circ\alpha:\pi_{1}(M)\to PSL_{2}(\mathbb{R})$. Hence, if $e(\nu_{M})=0$, then $\varphi$ lifts to a homomorphism $\pi_{1}(M)\to\widetilde{PSL}_{2}(\mathbb{R})\leq\mbox{Homeo}_{+}(\mathbb{R})$ [19, Lemma 2]. It follows that $\pi_{1}(M)$ is left-orderable [8, Theorem 1.1] and therefore, $M$ admits a horizontal foliation [8, Theorem 1.3]. ∎
@@ 5. Seifert fibred examples
We assume that $n\geq 3$. Then the base orbifold $\mathcal{B}$ of $M$ is spherical if $\chi(\mathcal{B})>0$, Euclidean if $\chi(\mathcal{B})=0$, and hyperbolic if $\chi(\mathcal{B})<0$. Orient $\mathcal{B}$ and endow $\nu_{M}$, the normal bundle to the Seifert fibres of $M$, with the induced orientation. The following lemma is key to our applications.
Lemma 5.1.
Suppose that $M$ is a Seifert fibred rational homology $3$-sphere. Then
@@
That any co-oriented taut foliation on $M$ must be horizontal is [7, Corollary 7] in the case $n=3$ and [1, Proposition 6.1] in general. Thus (1) holds. Consequently, the tangent bundle of any such foliation is isomorphic to $\nu_{M}$ (as an unoriented bundle), which implies (2). ∎
Theorem 1.8 follows immediately from this lemma and Theorem 1.11.
Proof of Theorem 1.7.
Let $K$ be the right-handed trefoil knot, and let $\mu$ and $\lambda$ denote the standard meridian and longitude of $K$. There is a natural Seifert fibre structure on the exterior $X=S^{3}\setminus\nu(K)$, whose regular fibre has slope $6\mu+\lambda$ on $\partial X$. Hence, $p/q$-surgery on $K$ yields a small Seifert fibred space whenever $p/q\neq 6$.
It is known that $K(p/q)$ admits co-orientable taut foliations when $p/q<1$ [25, §3] and hence has a left-orderable fundamental group [4]. On the other hand, assuming that $p>0$ and $p,q$ are coprime, [14, Theorem 1.7] shows that the Euler class of the co-orientable taut foliation constructed by Roberts in [25] vanishes if and only if
$|q|\equiv 1\pmod{p}$.
This completes the proof. ∎
@@
$(b,\frac{a_{1}}{b_{1}},\ldots,\frac{a_{n}}{b_{n}})\sim(b^{\prime},\frac{a_{1}^{\prime}}{b_{1}},\ldots,\frac{a_{n}^{\prime}}{b_{n}})$
if they differ by a finite sequence of moves each of which either multiplies the coordinates by 1-1 or is of the form described in Remark 3.1(1). Equivalent sets of Seifert invariants classify Seifert structures on a closed $3$-manifold up to homeomorphism. See Remark 3.1.
@@ 5.1. $\mathcal{B}$ is spherical
$(b,\frac{1}{2},\frac{2}{3},\frac{a_{3}}{b_{3}})\sim(b+1,\frac{1}{2},\frac{-1}{3},\frac{a_{3}}{b_{3}})\sim(-b-1,-\frac{1}{2},\frac{1}{3},-\frac{a_{3}}{b_{3}})\sim(-b-3,\frac{1}{2},\frac{1}{3},\frac{b_{3}-a_{3}}{b_{3}})$
@@ 5.2. $\mathcal{B}$ is Euclidean
Here, either $n=3$ and $(b_{1},b_{2},b_{3})$ is one of the triples $(2,3,6),(2,4,4)$, $(3,3,3)$ or $n=4$ and $(b_{1},b_{2},b_{3},b_{4})=(2,2,2,2)$. We can assume that $a_{i}=1$ in all cases. This is obvious when $b_{i}=2$ and using Lemma 5.2 and the arguments of §5.1 deals with the case that some $b_{i}>2$.
@@ 5.3. $\mathcal{B}$ is hyperbolic
We assume in this section that the Seifert invariants of
$M=M(0;b,\frac{a_{1}}{b_{1}},\ldots,\frac{a_{n}}{b_{n}})$
have been normalised so that $0<a_{i}<b_{i}$ for $1\leq i\leq n$.
@@
The following specific examples illustrate that there are many Seifert manifolds $M$ that have co-orientable taut foliations with non-zero Euler class.
Examples 5.3.
Using Lemma 5.2, if some $b_{i},b_{j}$ have a common factor greater than $2$, one can construct many examples $M$ with $e(\nu_{M})\neq 0$. On the other hand using the conditions (1), (2), (3) immediately above, one can arrange for many of these to have horizontal foliations.
In (1), (2), (3), and (4) below we give some other classes of examples of Seifert fibred manifolds which admit co-orientable taut foliations but none with vanishing Euler class. The proofs are all applications of Theorem 1.1; we omit the details.
(1) By Lemma 5.2, if $b_{1}=b_{2}=\cdots=b_{n}$ and $e(\nu_{M})=0$, then $a_{1}=a_{2}=\cdots=a_{n}$. The first set of examples deals with a special case of this situation.
For $n\geq 4$, consider the manifold $M=M(0;b,\frac{1}{c},...,\frac{1}{c})$. Then
  • (a)
    $M$ has a horizontal foliation if and only if either $2-n\leq b\leq-2$, or b=1b=-1 and $c\geq 3$.
  • (b)
    If $c\geq 2n-2$, then $e(\nu_{M})=0$ if and only if b=2nb=2-n.
The bound on $c$ in part (b) is optimal: if c=2n3c=2n-3 and b=1b=-1, then $e(\nu_{M})=0$.

### v3

@@ 3. The Euler class of the normal bundle of a Seifert fibration
Let $M$ be a closed, oriented Seifert fibred manifold whose base orbifold has an orientable underlying space. In this section, we compute the Euler class of $\nu_{M}$, the normal bundle to the Seifert fibres of $M$, and determine conditions for it to vanish.
@@ 3.1. Seifert notation
    This description of the Seifert fibred manifold is not unique. More precisely, $M\Bigl(g;\,\tfrac{a_{0}}{b_{0}},\ldots,\tfrac{a_{n}}{b_{n}}\Bigr)$ is homeomorphic to $M\Bigl(g;\,\tfrac{a^{\prime}_{0}}{b_{0}},\ldots,\tfrac{a^{\prime}_{n}}{b_{n}}\Bigr)$ by an orientation-preserving homeomorphism preserving fibres if and only if up to reordering,
@@
It follows from [17, Theorem 4] that the image of 1-1 under the transgression homomorphism $\delta$ is the Euler class $bu$ of the circle bundle. Hence $\mbox{kernel}(f^{*})$ is generated by $bu$. Then by (3.1.2), $e(\nu_{M})=0$ if and only $e(\tau_{B})=\pm\chi(B)u$ is a multiple of $bu$, which is condition (3.1.1).
@@
3.2. The triviality of the normal bundle of a Seifert fibring
Here is our main result.
Theorem 3.3.
  1. (1)
    $ma_{i}\equiv 1\,(\text{{\rm mod} }b_{i}),\text{ for }i=1,\cdots,n,\;\text{and}$
  2. (2)
    $me(M)=\chi(\mathcal{B})$.
Corollary 3.4.
Suppose that $M$ is a closed oriented Seifert fibred manifold $M$ such that the Euler class of $\nu_{M}$ vanishes. Then one of the following holds:
  1. (1)
    $M$ is an orientable $S^{1}$-bundle over the torus,
  2. (2)
    $e(M)=\chi(\mathcal{B})=0$,
  3. (3)
    $e(M)\neq 0,\,\chi(\mathcal{B})\neq 0,$ and
    $\frac{\chi(\mathcal{B})}{e(M)}\in\mathbb{Z}.$
The following lemma is key to applying these results to the Euler classes of foliations on Seifert fibred rational homology $3$-spheres.
Lemma 3.5.
Suppose that $M$ is a Seifert fibred rational homology $3$-sphere. Then
Proof.
That any co-oriented taut foliation on $M$ must be horizontal is [6, Corollary 7] in the case $n=3$ and [1, Proposition 6.1] in general. Thus (1) holds. Consequently, the tangent bundle of any such foliation is isomorphic to $\nu_{M}$ (as an unoriented bundle), which implies (2). ∎
Theorem 1.9 follows immediately from this lemma and Theorem 3.3.
@@ 3.2. The triviality of the normal bundle of a Seifert fibring
    From $\sigma$ one can define an obstruction $2$-cochain, denoted by $c_{\sigma}$, that represents the relative Euler class of the normal bundle $\nu_{M}$ in $H^{2}(M,\sqcup T_{i})$. By construction, the image of $c_{\sigma}]$ under $H^{2}(M,\sqcup T_{i})\to H^{2}(M)$ is the Euler class $e(\nu_{M})$. (See [18, Section 2] for instance.)
@@
4. Representations of Fuchsian groups and the normal bundle of a Seifert fibration
@@ 4.2. The normal bundle of the Seifert fibring of $M$ and discrete representations
Now proceed as in the proof of [2, Lemma A.3], starting from its third paragraph, to show that the $\mathbb{H}^{2}$-bundle $M\times_{\psi}\mathbb{H}^{2}\to M$ is topologically equivalent to the normal bundle $\nu_{M}$ and conclude, as in [2, Proposition A4], that $e(\nu_{M})=e(\varphi)=e(\bar{\varphi}\circ\alpha)$. Thus
@@
It is well-known that $e(\bar{\varphi})$, respectively $e(\varphi)$, is the obstruction to lifting the representation $\bar{\varphi}:\Gamma\to PSL_{2}(\mathbb{R})$, respectively $\varphi:\pi_{1}(M)\to PSL_{2}(\mathbb{R})$, to $\widetilde{PSL}_{2}(\mathbb{R})$ [31, Lemma 2].
@@
(4.2.4) $\theta(e(E_{M}))=-l\left(b+\sum_{i=1}^{n}\frac{a_{i}}{b_{i}}\right)=-le(M),$
@@
(4.2.5) $\theta(e(\bar{\varphi}))=l\left(n-\chi(|\mathcal{B}|)-\sum_{i}\frac{1}{b_{i}}\right)=-l\chi(\mathcal{B})$
@@
We saw in (4.2.1) that that $e(\nu_{M})=\alpha^{*}(e(\bar{\varphi}))=e(\bar{\varphi}\circ\alpha)=e(\varphi)$, where $\varphi=\bar{\varphi}\circ\alpha:\pi_{1}(M)\to PSL_{2}(\mathbb{R})$. Hence, if $e(\nu_{M})=0$, then $\varphi$ lifts to a homomorphism $\pi_{1}(M)\to\widetilde{PSL}_{2}(\mathbb{R})\leq\mbox{Homeo}_{+}(\mathbb{R})$ [31, Lemma 2]. It follows that $\pi_{1}(M)$ is left-orderable [7, Theorem 1.1] and therefore, $M$ admits a horizontal foliation [7, Theorem 1.3]. ∎
@@ 5. Seifert fibred examples
We assume that $n\geq 3$. Then the base orbifold $\mathcal{B}$ of $M$ is spherical if $\chi(\mathcal{B})>0$, Euclidean if $\chi(\mathcal{B})=0$, and hyperbolic if $\chi(\mathcal{B})<0$. Orient $\mathcal{B}$ and endow $\nu_{M}$, the normal bundle to the Seifert fibres of $M$, with the induced orientation.
Proposition 5.1.
Let $M$ be a Seifert fibred rational homology sphere with $n$ exceptional fibres. If $n\geq|H_{1}(M)|+3$, then $M$ admits a coorientable taut foliation.
@@
Suppose that $M$ can be expressed in the form
$M=M\left(0;b,\frac{a_{1}}{b_{1}},\dots,\frac{a_{n}}{b_{n}}\right),$
where $0<\frac{a_{1}}{b_{1}}<1$ for all $i$. We give two proofs, one from the viewpoint of Heegard Floer homology, the other from the viewpoint of foliations transverse to the Seifert fibres of $M$.
The Heegard Floer approach. Up to reversing orientation, we can assume that $e(M)<0$.
Suppose that $M$ does not admit a coorientable taut foliation, or equivalently, that $M$ is an $L$-space (Lemma 3.5 and [30]). Némethi [33, § 11] provides an algorithm for calculating the Heegaard Floer homology of $M$ in terms of its Seifert invariants. This algorithm shows that if $M$ is an $L$-space then the function
$\Delta(i)=1-ib-\sum_{j=1}^{n}\left\lceil\frac{ia_{j}}{b_{j}}\right\rceil$
satisfies $\Delta(i)\geq 0$ for all $i\geq 0$. [^4] Actually this is equivalent to being an $L$-space. In particular, one calculates that
$\Delta(1)=1-b-n\geq 0$
and that
$\Delta((b_{1}\cdots b_{n})-1)=1+|H_{1}(M)|+b\geq 0.$
Summing these two implies that $n\leq|H_{1}(M)|+2$, as required.
The transverse foliation approach. After reversing orientation we have $e(M)>0$, and after reindexing we can suppose that $b_{1}\geq b_{i}$ for each $i$. Then with $p=b_{1}b_{2}\cdots b_{n}$ and $\sigma=\sum_{i=1}^{n}a_{i}/b_{i}$, we have $p(b+\sigma)=pe(Y)=|H_{1}(M)|$. Therefore, as $|H_{1}(M)|\leq n-3<nb_{2}b_{3}\cdots b_{n}$,
$\frac{|H_{1}(M)|}{p}-n<\frac{|H_{1}(M)|}{p}-\sigma=b\leq\frac{|H_{1}(M)|}{p}-\frac{n}{b_{1}}=\frac{1}{b_{1}}\left(\frac{|H_{1}(M)|-nb_{2}b_{3}\cdots b_{n}}{b_{2}b_{3}\cdots b_{n}}\right)<0.$
It follows that $1-n\leq b\leq-1$.
If $2-n\leq b\leq-2$, $M$ admits a co-orientable taut foliation by [23, Theorem 2].
If b=1b=-1, set $m=p+|H_{1}(M)|-(n-2)$. Then $2\leq m<p$ and if $p_{i}=p/b_{i}\in\mathbb{Z}$ we have
  • $a_{1}p_{1}+a_{2}p_{2}+\cdots+a_{n}p_{n}=p\sigma=p+|H_{1}(M)|=m+n-2$;
  • $a_{i}/b_{i}=a_{i}p_{i}/p<a_{i}p_{i}/m$.
It follows that the $n$-tuple $(a_{1}/b_{1},\ldots,a_{n}/b_{n})$ lies in $R_{n}(m)$, where
$R_{n}(m)=\{(\gamma_{1},\gamma_{2},\ldots,\gamma_{n})\in(0,1)^{n}\;|\;\exists\;(r_{1},r_{2},\ldots,r_{n})\in\mathbb{Z}^{n}\mbox{ such that }$
[…]
Corollary 3.3 of [23] now implies that $M$ admits a co-orientable taut foliation.
A similar argument can be applied when b=1nb=1-n, after observing that
$M=M\left(0;1-n,\frac{a_{1}}{b_{1}},\dots,\frac{a_{n}}{b_{n}}\right)$ $\cong$ $M\left(0;n-1,-\frac{a_{1}}{b_{1}},\dots,-\frac{a_{n}}{b_{n}}\right)$
$=$ $M\left(0;-1,\frac{b_{1}-a_{1}}{b_{1}},\ldots,\frac{b_{n}-a_{n}}{b_{n}}\right)$
@@
Proof of Theorem 1.13.
Let $G$ be a finite abelian group of containing an element of order at least $3$. We may write $G\cong\frac{\mathbb{Z}}{d_{1}}\oplus\dots\oplus\frac{\mathbb{Z}}{d_{k}}$, where $d_{i+1}$ divides $d_{i}$ for each $i$ and $d_{1}\geq 3$. Consider the Seifert fibred space
(5.0.1) $M=M\left(0;0,\frac{a}{d_{1}^{2}},\frac{1}{d_{1}},\dots,\frac{1}{d_{k}}\right),$
where $a=-1-\sum_{i=1}^{k}\frac{d_{1}^{2}}{d_{i}}$. For such an $a$ we have $a\equiv-1\;(\bmod\;d_{1})$. Consequently $a$ is coprime to $d_{1}^{2}$. Moreover, as $d_{1}>2$, we see from Lemma 5.2 that $e(\nu_{M})\neq 0$. Calculating the homology using (4.0.1) shows that $H_{1}(M)\cong G$ (compare [19, § 4]).
Next observe that we may perform surgery on the regular fibres in order to obtain a new Seifert fibered space with the same first homology group.
More precisely, suppose that $Y$ is the Seifert fibred rational homology $3$-sphere
$Y=M\left(0;b,\frac{a_{1}}{b_{1}},\dots,\frac{a_{n}}{b_{n}}\right)$
oriented so that $e(Y)>0$ and let $H=|H_{1}(Y)|$ denote the order of the first homology of $Y$. Then one can verify that the manifold
$Y^{\prime}=M\left(0;b,\frac{a_{1}}{b_{1}},\dots,\frac{a_{n}}{b_{n}},\frac{-H^{2}}{(b_{1}\cdots b_{n})H+1}\right),$
satisfies $H_{1}(Y)\cong H_{1}(Y^{\prime})$. First, one can calculate that
$|H_{1}(Y^{\prime})|=(b_{1}b_{2}\cdots b_{n})(H(b_{1}b_{2}\cdots b_{n})+1)e(Y^{\prime})=|H_{1}(Y)|.$
On the other hand, consideration of the presentations of $H_{1}(Y)$ and $H_{1}(Y^{\prime})$ obtained from (4.0.1) shows that there is an epimorphism $H_{1}(Y^{\prime})\to H_{1}(Y)$ induced by sending $x_{i}$ to $x_{i}$ for $1\leq i\leq n$, $x_{n+1}$ to $1$, and $h$ to $h$. Thus $H_{1}(Y^{\prime})\cong H_{1}(Y)$.
Now starting with the manifold $M$ as defined in (5.0.1), and applying the above construction, we obtain an infinite sequence of Seifert fibred spaces, none of which admit a co-orientable taut foliation with Euler class zero by condition (1) of Theorem 3.3. It follows from Proposition 5.1 that if
$Y=M\left(0;b,\frac{a_{1}}{b_{1}},\dots,\frac{a_{n}}{b_{n}}\right)$
is a Seifert fibred rational homology sphere which is an $L$-space or, equivalently, does not admit a coorientable taut foliation, then $n\leq 2+|H_{1}(Y)|$. Thus, the manifolds in this sequence will eventually admit co-orientable taut foliations. ∎
@@
$\left(b,\frac{a_{1}}{b_{1}},\ldots,\frac{a_{n}}{b_{n}}\right)\sim\left(b^{\prime},\frac{a_{1}^{\prime}}{b_{1}},\ldots,\frac{a_{n}^{\prime}}{b_{n}}\right)$
if they differ by a finite sequence of moves each of which either multiplies the coordinates by 1-1 or is of the form described in Remark 3.1(1). Equivalence classes of Seifert invariants classify Seifert structures on a closed $3$-manifold up to homeomorphism. See Remark 3.1.
@@ 5.1. $\mathcal{B}$ is spherical
$\left(b,\frac{1}{2},\frac{2}{3},\frac{a_{3}}{b_{3}}\right)\sim\left(b+1,\frac{1}{2},\frac{-1}{3},\frac{a_{3}}{b_{3}}\right)\sim\left(-b-1,-\frac{1}{2},\frac{1}{3},-\frac{a_{3}}{b_{3}}\right)\sim\left(-b-3,\frac{1}{2},\frac{1}{3},\frac{b_{3}-a_{3}}{b_{3}}\right)$
@@
Remark 5.3.
Let $T(r,s)$ be the $(r,s)$-torus knot, where $r<s$ are relatively prime integers greater than 1, and let $g$ denote its genus. If $p/q\neq rs$ then $M=T(r,s)(p/q)$ has a Seifert fibration induced from that on the exterior of $T(r,s)$, and we can use Theorem 3.3 to determine whether or not the Euler class $e(\nu_{M})$ is zero. It turns out that if $(r,s)\neq(2,3)$, then $e(\nu_{M})=0$ if and only if $(2g-1)|q|\equiv 1(\text{{\rm mod} }p)$. Since $g>1$, this implies that p/q<2g1p/q<2g-1, and therefore $M$ admits a co-orientable taut foliation. On the other hand, for the trefoil $T(2,3)$ we get that $e(\nu_{M})=0$ if and only if $|q|\equiv 1(\text{{\rm mod} }p)$ and $p/q<6$. Thus either $p/q<1$, in which case $M$ has a co-orientable taut foliation, or $p/q=1,2,3,4$ or $5$. The latter five manifolds $M$ are, respectively, those with Seifert invariants (1,1/2,1/3,1/5),(1,1/2,1/3,1/4),(1,1/2,1/3,1/3),(1,1/2,1/2,1/3)(-1,1/2,1/3,1/5),(-1,1/2,1/3,1/4),(-1,1/2,1/3,1/3),(-1,1/2,1/2,1/3), and $L(5,1)$. As these have finite fundamental groups, they do not admit co-orientable taut foliations.
@@ 5.2. $\mathcal{B}$ is Euclidean
Here, either $n=3$ and $(b_{1},b_{2},b_{3})$ is one of the triples $(2,3,6),(2,4,4)$, $(3,3,3)$ or $n=4$ and $(b_{1},b_{2},b_{3},b_{4})=(2,2,2,2)$. We can assume that $a_{i}=1$ in all cases. This is obvious when $b_{i}=2$ and using Lemma 5.2 and the arguments of §5.1 deal with the case that some $b_{i}>2$.
@@ 5.3. $\mathcal{B}$ is hyperbolic
Throughout this subsection, we assume that $\mathcal{B}=S^{2}(b_{1},...,b_{n})$ is hyperbolic. We also assume that $M=M(0;b,\frac{a_{1}}{b_{1}},\ldots,\frac{a_{n}}{b_{n}})$ is a Seifert manifold with base orbifold $\mathcal{B}$ whose Seifert invariants have been normalised so that $0<a_{i}<b_{i}$ for $1\leq i\leq n$.
Remark 5.4.
@@
Theorem 5.5.
Let $\mathcal{B}$ be as above.
  1. (1)
    There exists a Seifert fibred rational homology sphere $M$ with base orbifold $\mathcal{B}$ such that $M$ has a horizontal foliation and $e(\nu_{M})=0$.
  2. (2)
    With exactly 58 exceptions $\mathcal{B}$, there exists a Seifert fibred rational homology sphere $M$ with base orbifold $\mathcal{B}$ such that $M$ has a horizontal foliation and $e(\nu_{M})\neq 0$.
@@
Thus the manifolds $M$ in part (2) have co-orientable taut foliations but none with Euler class zero.
To prove part (1) we take $M=M(0;2-n,1/b_{1},...,1/b_{n})$. It follows from Remark 5.4 that $M$ has a horizontal foliation. (When $n=3$, use the fact that $\mathcal{B}$ is hyperbolic.) Taking $m=1$ in Theorem 3.3 shows that $e(\nu_{M})=0$.
The exceptions in part (2) are explicitly listed in the Appendix. We will not include a complete proof of part (2); we content ourselves with proving two partial results, Theorems 5.6 and 5.7 below, to illustrate the general strategy.
First we introduce some notation. Let $X=-\chi(\mathcal{B})=(n-2)-\sum_{i=1}^{n}{1/b_{i}}$. Since $\mathcal{B}$ is hyperbolic, $X>0$. Next let $M$ be a Seifert fibre space with base orbifold $\mathcal{B}$; thus $M$ is of the form $M(0;b,a_{1}/b_{1},...,a_{n}/b_{n})$. Then
$e(M)=-b-\sum_{i=1}^{n}{a_{i}/b_{i}}=X-\left((n-2)+b+\sum_{i=1}^{n}\frac{a_{i}-1}{b_{i}}\right)$
Let $\alpha=(n-2)+b+\sum_{i=1}^{n}{(a_{i}-1)/b_{i}}$. Then
$e(M)=X-\alpha$
By Theorem 3.3, $e(\nu_{M})=0$ if and only if there exists an integer $m$ such that $ma_{i}\equiv 1$ (mod $b_{i}$) for $1\leq i\leq n$ and $X=m(X-\alpha)$, i.e. $m\alpha=(m-1)X$. We note that if $X-\alpha>0$ then $M$ is a rational homology sphere, and if in addition $\alpha\neq 0$ then $m>1$.
Theorem 5.6.
Let $\mathcal{B}=S^{2}(b_{1},...,b_{n})$ be as above, where $n\geq 7$. If $n=7$, assume that some $b_{i}$ is greater than $2$. Then there is a Seifert fibred rational homology sphere $M$ with base orbifold $\mathcal{B}$ such that $M$ has a horizontal foliation and $e(\nu_{M})\neq 0$.
Proof.
We take $M$ to be $M(0;3-n,1/b_{1},...,1/b_{n})$. Then $M$ has a horizontal foliation by Remark 5.4.
We have $X\geq(n-2)-n/2=(n-4)/2$, with strict inequality if $n=7$. Hence $X>3/2$. Also, $\alpha=1$, so $M$ is a rational homology sphere.
If $e(\nu_{M})=0$ then there exists an integer $m>1$ such that $m=m\alpha=(m-1)X>(3/2)(m-1)$. Therefore $m<3$, so $m=2$. But then $m\not\equiv 1$ (mod $b_{i}$) for any $i$. ∎
The exclusion of the case $\mathcal{B}=S^{2}(2,2,...,2)$, $n=7$, is necessary. To see this, note that here X=(72)7/2=3/2X=(7-2)-7/2=3/2. The manifolds $M$ with base orbifold $\mathcal{B}$ are those of the form M(0;b,1/2,1/2,,1/2)M(0;b,1/2,1/2,...,1/2). By Remark 5.4 $M$ has a horizontal foliation if and only if $5\leq b\leq-2$, i.e. $\alpha=k$, $0\leq k\leq 3$. The corresponding manifold $M_{k}$, say, has $e(\nu_{M_{k}})=0$ if and only if there exists an odd integer $m_{k}$ such that $X=m_{k}(X-k)$. Hence $m_{k}=3/(3-2k)$, so $m_{0}=1,m_{1}=3,m_{2}=-3$, and $m_{3}=-1.$ (In fact $M_{3-k}\cong-M_{k}.)$ Thus every $M$ with base orbifold $\mathcal{B}$ that has a horizontal foliation has $e(\nu_{M})=0$, so $\mathcal{B}(2,2,...,2)$, $n=7$, is one of the exceptions in part (2) of Theorem 5.5.
The next theorem deals with the case $n=3$ when each $b_{i}$ is at least $5$.
Theorem 5.7.
Let $\mathcal{B}=S^{2}(b_{1},b_{2},b_{3})$, where $b_{i}\geq 5$, $1\leq i\leq 3$, and the $b_{i}$ are not all $6$. Then there is a Seifert fibred rational homology sphere $M$ with base orbifold $\mathcal{B}$ such that $M$ has a horizontal foliation and $e(\nu_{M})\neq 0$.
The proof will use the following lemma.
Lemma 5.8.
If either $c=5$ or $c\geq 7$ then there exists a prime $p$ which does not divide $c$ such that $(p-1)/c\leq 1/3$ and $p/c<1/2$.
Proof.
If $c$ is odd take $p=2$.
If $c$ is even and not a multiple of $3$, take $p=3$.
So we may suppose that $c=6n$, where $n>1$. By Bertrand’s Postulate there exists a prime $p$ such that $n+2<p<2n+2$. The left-hand inequality implies that $p$ does not divide $c$ and the right-hand inequality shows that $p-1\leq 2n$. Therefore $(p-1)/c\leq 2n/6n=1/3$. Also $p/c\leq 1/3+1/c<1/2$. ∎
Proof of Theorem 5.7.
If two of the $b_{i}$, say $b_{1}$ and $b_{2}$, are equal to $5$ then we take $M=M(0;-1,1/5,2/5,1/b_{3})$; see Lemma 5.2. We may therefore assume that X>13/5=2/5.X>1-3/5=2/5.
By hypothesis, some $b_{i}$, say $b_{1}$, is not equal to $6$. Then by Lemma 5.8 there is a prime $p$ such that $(p-1)/b_{1}\leq 1/3$ and $p/b_{1}<1/2$. Let $M=M(0;-1,p/b_{1},1/b_{2},1/b_{3})$. Since $p/b_{1}<1/2$, $M$ has a horizontal foliation by Remark 5.4. Also, $\alpha=(p-1)/b_{1}$, so $X-\alpha>0$. Hence $M$ is a rational homology sphere. If $e(\nu_{M})=0$ then there exists $m>1$ such that $m\alpha=(m-1)X$. Therefore m/3>(2/5)(m1)m/3>(2/5)(m-1). Hence 5m>6(m1)5m>6(m-1), so $m<6$. But then, since $b_{2}\geq 5$, $m\not\equiv 1$ (mod $b_{2}$). ∎
It is necessary to exclude the case $\mathcal{B}=S^{2}(6,6,6)$. For it is easy to see from Remark 5.4 that the only manifold $M$ with base orbifold $\mathcal{B}$ that has a horizontal foliation is (up to orientation) M(0;1,1/6,1/6,1/6)M(0;-1,1/6,1/6,1/6), which has $e(\nu_{M})=0$.
Appendix A Exceptions to Theorem 5.5
Here we list the orbifolds $S^{2}(b_{1},...,b_{n})$ that are exceptions in part (2) of Theorem 5.5. We adopt the notational convention stated at the beginning of §5.3, and also assume that $b_{1}\leq\cdots\leq b_{n}$.
Let $M=M(0;b,a_{1}/b_{1},...,a_{n}/b_{n})$ be a Seifert fibre space with base orbifold $S^{2}(b_{1},...,b_{n})$. We distinguish three types of exceptional $S^{2}(b_{1},...,b_{n})$.
  1. (1)